# Jonathan Roberts1, Kai Han2, Samuel Albanie

### v2

@@ 4.3 Overall results
Claude 3 Haiku [3] 14.2 6.6 8.8 3.9 9.0
Claude 3 Sonnet [3] 15.3 8.6 4.5 4.8 9.2
Claude 3.5 Sonnet [4] 41.8 15.5 11.0 10.0 21.7
GPT-4 Turbo [27] 18.5 8.5 4.9 3.5 10.0
GPT-4o mini [29] 15.8 6.8 5.7 2.9 8.7
GPT-4o [28] 24.7 10.8 9.2 3.5 13.6
Gemini 1.0 Pro Vision [10] 20.2 5.8 6.9 6.1 10.5
Gemini 1.5 Flash [37] 28.5 11.5 8.4 9.0 15.6
Gemini 1.5 Pro [37] 34.2 11.4 13.3 6.5 18.1
Reka Edge [34] 11.8 8.7 11.6 1.9 9.4
Reka Flash [34] 13.2 10.1 6.3 3.9 9.3
Reka Core [34] 1.7 0.0 4.3 0.3 1.5
@@
CogVLM-Chat [43] 7.0 4.9 5.1 3.9 5.4
Qwen-VL-Chat [5] 10.2 6.6 5.1 2.9 6.8
OmniLMM-3b [32] 6.7 4.9 4.1 4.5 5.2
TransCore-M [36] 7.9 9.2 7.6 3.9 7.6
Yi-VL-6b [1] 5.6 8.6 7.1 4.2 6.7
Yi-VL-34b [1] 7.6 5.9 5.5 2.3 5.8
LLaVA-1.5 7b [18] 4.7 7.5 6.5 4.8 6.0
LLaVA-1.5 13b [18] 5.0 7.7 8.4 3.9 6.5
Table 3: Overall results on the GRAB tasks. The highest open and closed-source model scores for each task are highlighted and bold.
@@
Table 4: GRAB per-category performance across all tasks. The highest open and closed-source model scores for each category are highlighted and bold.
We present the overall results on GRAB in Tab. 3. A clear observation is that our benchmark proves challenging for frontier models, with low overall performances from all models (see also Fig. 1). The highest performing model, Claude 3.5 Sonnet, achieves a score of just 21.7%. As expected, the highest scores are attained on the Properties task, which includes questions that just require the derivation of a single graph property. Lower scores are achieved on the Functions and Series tasks, likely due to the added complexity of computing the mean of a given property over multiple functions/series. There is a macro-trend of slightly higher accuracy on the functions task compared to the series task, possibly due to the irregularity introduced by the noise in the data. Significantly lower accuracy is attained on the Transforms task. A contributing factor to these scores is the degree to which the models follow output format instructions. Given our exact string matching evaluation protocol, any deviation from the expected output is deemed incorrect. Despite potentially having superior mathematical reasoning abilities, models that struggle to precisely follow instructions or have proclivities to generate verbose answers suffer severe performance degradation. With the exception of Reka Core (which showed poor output format instruction following capabilities), all closed-source models outperformed the leading open-source model. However, the open-source models did consistently generate answers in the correct format.
In this low-scoring regime where even the best models struggle, a comparison of model performance should be made with the caveat that all models can barely perform the tasks in the benchmark. We anticipate our benchmark will yield most benefit when evaluating the next generation of LMMs.
4.4 Category analysis
A decomposition of model performance into question category is provided in Tab. 4. Significant variance is shown across categories, with some models much better suited for certain categories than others. Overall, the highest accuracies are attained on the Counting category, which encompasses simple questions determining the number of series shown in each plot, or the approximate number of data points. On the other hand, no model correctly answered a single question from the Functions category. This challenging category requires correct derivation of multiple parameters (e.g. the gradient and intercept for linear functions) and correct formatting of the output. Inspection of the model output suggests adherence to the strict output format was not an issue: the models were unable to correctly determine the function parameters. However, in a number of cases, proposed functions were close to the correct answers, e.g., with all parameters correct apart from one that was off by one. Minor perturbations to the graph appearance did result in a very small number of correct answers.
@@ 4.5 Evaluation protocols
Figure 5: GRAB accuracy with different automatic evaluation protocols. Performance scores are shown for each model for (i) exact matching, (ii) answer extraction with an open-source LLM then exact matching, and (iii) answer extraction with a closed-source LLM then exact matching. Accuracy differences relative to the exact matching protocol are annotated.
@@
In addition to the exact matching evaluation protocol (see Sec. 4.2.3), we explore model performance scores with more lenient automatic evaluation protocols. Specifically, we utilised an LLM to parse the generated output from the LMMs and extract the answer text, before evaluating it via exact matching as before. We carried out this evaluation method using both a frontier LLM (Gemini 1.5 Pro) and a small open-source LLM (LLaMA 3.1 8b-instruct [24]). Details of the prompts used can be found in the Appendix. Based on random sampling, the LLMs could extract specific answers from verbose output with high accuracy, however, performed poorly when provided the correct answer and prompted to actually determine the correctness of a given answer. A comparison of the different evaluation protocols can be found in Fig. 5. On the whole, there is almost no difference between the exact matching and automatic evaluation protocols. Similarly, there is little difference between extraction with a frontier closed-source and a small (8B parameter) open-source model, suggesting this approach can feasibly be conducted with little compute. However, we advocate for the adoption of exact matching without the use of an LLM as the standard evaluation approach going forwards as most models are clearly able to sufficiently follow output instructions and it provides a better signal of model capabilities. Cases where there are differences between the performance protocols indicate models that are the poorest instruction followers.
4.6 Complexity analysis
The Functions, Series and Transforms tasks were designed to include questions spanning a controllable complexity scale. For the Functions and Series tasks, complexity refers to the number of additional functions or series present on the graph, while for the Transforms task, complexity represents the number of transformations applied to the function. In both cases, the complexity score ranges from 0 (graphs with a single function or series and no transformations) to 9. In Fig. 6, we present performance decomposed along this complexity scale for a representative selection of models. Intuitively, we expect performance to decrease as complexity increases due to the added challenge of determining the quantity of interest from additional functions/series or after additional transforms. However, the complexity plots for the evaluated models follow two distinct patterns. For the better performing models, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, performance does clearly decrease as complexity increases from 0 to 3, before remaining around 10%. For the weaker models, however, the results fluctuate around 10% across the entire complexity domain. In these cases, even the lowest complexity questions are too challenging. As expected, the questions requiring decimal precision proved much harder than integer precision questions.
4.7 Question format
@@
5 Conclusions
We introduce GRAB, a GRaph Analysis Benchmark for LMMs. We leverage synthetic data generation to create high-quality, difficult questions that prove challenging for frontier LMMs, with the best-performing model attaining a score of 21.7%. On GRAB, we evaluate a suite of 20 closed and open-source models, attaining a broad profile of the abilities of the current generation of LMMs. During inference, we prompt the models using lightweight prompts with clear output format instructions and assess answer correctness by comparing the output string with the ground truth. This approach jointly assesses the abilities of models to perform the GRAB reasoning tasks and reliably follow output instructions, enabling straightforward evaluation without the need for human or automatic evaluation using LLMs. A consequence of this harsh evaluation protocol is that some models commonly regarded as ‘frontier’ – that are prone to verbose generation – perform relatively poorly. We directly compare performance results using exact matching evaluation with automatic LLM evaluation and find little difference, suggesting most current models are able to sufficiently follow output instructions, and therefore validating exact matching as a suitable evaluation protocol. Additionally, we conduct a series of studies revealing the categories the models struggle with and the impact of different question types. We release our dataset and evaluation codebase for the community to use and hope our work encourages research towards the next generations of LMMs.
Acknowledgements
This work was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1), an Isaac Newton Trust grant, a research gift from Google, an EPSRC HPC grant, the Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27208022), and HKU Seed Fund for Basic Research. Samuel would like to acknowledge the support of Z. Novak and N. Novak in enabling his contribution.
Appendix
We structure this Appendix into 4 parts. In (1), we include examples of the prompts used for inference and LLM evaluation. Then, in (2) we include details of the compute costs of our work. Next, in (3) we present results on an OCR-based graph analysis question set. Finally, in (4) we detail the specific versions of the API models used in this work, aiding reproducibility.
Appendix A Prompts
@@
Appendix B Compute
As we restricted the size of our benchmark to just 2170, the total compute required for this work is relatively low. Negligible compute was needed to create the synthetic images. Inference with the closed-source models was carried out via API calls. Using a single NVIDIA A100-80GB GPU, inference on the entire GRAB benchmark using LLaVA-1.5 7b can be carried out in approximately 30 minutes, using a single process (inference time is significantly reduced with multiprocessing).
Appendix C OCR experiments
Model Accuracy (%)
Random chance 20.0
Closed-source LMMs
Claude 3 Haiku 42.9
Claude 3 Sonnet 91.4
Gemini 1.0 Pro Vision 100.0
Gemini 1.5 Flash 100.0
Gemini 1.5 Pro 100.0
GPT-4 Turbo 97.1
GPT-4V 97.1
GPT-4o 82.9
Reka Core 97.1
Reka Edge 88.6
Reka Flash 100.0
Open-source LMMs
TransCore-M 80.0
Yi-VL-6b 51.4
OmniLMM-3b 65.7
Qwen-VL-Chat 71.4
Table 6: Accuracy scores on OCR experiments.
Appendix D API model versions
@@
  • Gemini-Pro:
    gemini-1.0-pro-001

### v3

@@ 4.4 Real splits analysis
Claude 3 Haiku [3] 9.8 4.9 9.5 9.3 9.2
Claude 3 Sonnet [3] 9.8 14.6 12.8 12.0 12.4
Claude 3.5 Sonnet [4] 14.6 17.0 18.6 21.1 19.6
GPT-4 Turbo [36] 9.8 4.9 6.6 8.3 7.5
GPT-4o mini [38] 7.3 2.4 4.7 3.3 4.0
GPT-4o [37] 29.3 9.8 18.0 16.3 17.3
Gemini 1.5 Flash [49] 31.7 17.1 16.3 16.7 17.1
Gemini 1.5 Pro [49] 34.4 24.4 21.9 17.2 20.3
Reka Flash [46] 19.5 7.3 7.9 11.4 10.0
Reka Core [46] 4.9 0.0 1.0 1.4 1.3
@@
CogVLM-Chat [55] 14.6 7.3 11.0 9.9 10.5
Qwen-VL-Chat [8] 12.2 4.9 2.9 5.6 4.6
OmniLMM-3b [44] 7.3 2.4 6.6 6.0 6.2
TransCore-M [48] 2.4 9.8 7.6 9.1 8.2
Yi-VL-6b [1] 7.3 4.9 9.9 10.1 9.7
Yi-VL-34b [1] 4.9 4.9 7.8 7.8 7.5
LLaVA-1.5 7b [27] 4.9 4.9 9.5 8.1 8.5
LLaVA-1.5 13b [27] 7.3 4.9 8.9 9.3 8.9
Gemini 2.0 Flash [15] 22.0 17.0 18.8 20.5 19.6
Gemini 2.5 Flash [17] 36.6 22.0 29.5 29.3 29.4
Table 4: Performance on the GRAB ‘Real’ task splits. Overall scores are weighted averages based on task size.
4.5 Category analysis
@@
Table 5: GRAB per-category performance. The highest open and closed-source model scores for each category are highlighted and bold. To isolate per-category performance and avoid the effects of noise and distractors, the results in this table do not include the Real task.
A decomposition of model performance into question category is provided in Tab. 5. Significant variance is shown across categories, with some models much better suited for certain categories than others. Overall, the highest accuracies are attained on the Counting category, which encompasses simple questions determining the number of series shown in each plot, or the approximate number of data points. On the other hand, no model correctly answered a single question from the Functions category. This challenging category requires correct derivation of multiple parameters (e.g. the gradient and intercept for linear functions) and correct formatting of the output. Inspection of the model output suggests adherence to the strict output format was not an issue: the models were unable to correctly determine the function parameters. However, in a number of cases, proposed functions were close to the correct answers, e.g., with all parameters correct apart from one that was off by one. Minor perturbations to the graph appearance did result in a small number of correct answers.
4.6 Complexity analysis
The Functions, Series and Transforms tasks were designed to include questions spanning a controllable complexity scale. For Functions and Series, complexity refers to the number of additional functions or series present on the graph, while for Transforms, it represents the number of transformations applied to the function. In both cases, the complexity score ranges from 0 (single function/series and no transformations) to 9. In Fig. 5, we present performance decomposed along this complexity scale for a representative selection of models. Intuitively, we expect performance to decrease as complexity increases due to the added challenge of determining the quantity of interest from additional functions/series or after additional transforms. However, the complexity plots for the evaluated models follow two distinct patterns. For the better performing models, such as Claude 3.5 Sonnet and Gemini 1.5 Pro, performance does clearly decrease as complexity increases from 0 to 3, before remaining around 10%. For the weaker models, however, the results fluctuate around 10% across the entire complexity domain. In these cases, even the lowest complexity questions are too challenging. As expected, the questions requiring decimal precision proved much harder than integer precision questions.
@@ 4.7 Evaluation protocols
Figure 6: Overall accuracy on the Properties, Functions, Series and Transforms tasks with different automatic evaluation protocols. Scores are shown for exact matching and answer extraction (with different LLMs) and then exact matching. All accuracy differences relative to the exact matching protocol are annotated – in nearly every case, there is either minor or no difference in score.
@@
In addition to the exact matching evaluation protocol (§4.2.3), we explore model performance with more lenient automatic evaluation protocols. Specifically, we utilised an LLM to parse the generated output from the LMMs and extract the answer text, before evaluating it via exact matching as before. We carried this out using both a frontier LLM (Gemini 1.5 Pro) and a small open-source LLM (LLaMA 3.1 8b-instruct [33]). Details of the prompts used can be found in the Appendix. We found the LLMs could extract specific answers from verbose output with high accuracy but performed at determining the correctness of a model response given the ground truth answer. A comparison of the different evaluation protocols can be found in Fig. 6. On the whole, there is almost no difference between the exact matching and automatic evaluation protocols. Similarly, there is little difference between extraction with a frontier closed-source and a small open-source model, suggesting this approach can feasibly be conducted with little compute. However, we advocate for the adoption of exact matching without the use of an LLM as the standard evaluation approach going forwards as most models are clearly able to sufficiently follow output instructions and it provides a better signal of model capabilities. Cases where there are differences between the performance protocols indicate models that are the poorest instruction followers.
4.8 Additional ablations
Additional ablations and analyses can be found in the Appendix, including a comparison of the Matplotlib and Seaborn plotting libraries, alternative evaluation metrics, results on an OCR-based question set, and an analysis of performance with different question formats.
4.9 GRAB-Lite
Accuracy (%)
Properties Functions Series Transforms Real Overall
Model (100) (100) (100) (100) (100) (500)
GPT-4o [37] 21.0 7.0 10.0 6.0 19.0 12.6
Claude 3.5 Sonnet [4] 39.0 15.0 11.0 13.0 20.0 19.6
Gemini 1.5 Pro [49] 23.0 10.0 14.0 7.0 25.0 15.8
GPT-5 [43] 59.0 34.0 33.0 63.0 42.0 46.2
GPT-5 mini [43] 55.0 40.0 32.0 56.0 33.0 43.2
GPT-5 nano [43] 36.0 34.0 29.0 33.0 19.0 30.2
GPT-4.1 [42] 31.0 21.0 30.0 29.0 24.0 27.0
o1 [39] 27.0 15.0 28.0 26.0 25.0 24.2
Claude Sonnet 4.5 [7] 47.0 34.0 39.0 48.0 29.0 39.4
Claude Sonnet 4 [6] 37.0 31.0 31.0 29.0 24.0 30.4
Claude 3.7 Sonnet [5] 36.0 13.0 13.0 11.0 10.0 16.6
Gemini 2.5 Pro [19] 54.0 43.0 31.0 55.0 38.0 44.2
Gemini 2.5 Flash [17] 34.0 27.0 29.0 22.0 30.0 28.4
Gemini 2.5 Flash Lite [18] 18.0 5.0 11.0 18.0 10.0 12.4
Gemini 2.0 Flash [15] 41.0 25.0 18.0 37.0 28.0 29.8
Gemini 2.0 Flash Lite [16] 14.0 13.0 9.0 14.0 12.0 12.4
Grok 4 [58] 23.0 15.0 28.0 22.0 20.0 21.6
Table 6: Results on GRAB-Lite, a lightweight 500-question subset of GRAB designed to evaluate costly reasoning models.
Recently, there has been a shift towards reasoning models that leverage inference-time scaling and spend more effort “thinking” [39]. While this strategy can lead to better-reasoned output, it requires more output tokens, which increases the cost of inference significantly and constrains the scale of evaluation. This creates a pressing need for lightweight benchmarks. To address this, we introduce a small 500-question version of our GRAB, called GRAB-Lite. GRAB-Lite is an approximately category, complexity, and task-balanced composition of GRAB with synthetic graphs generated as an equal split between Matplotlib and Seaborn. We evaluate 3 strong GRAB baselines along with newer frontier models (many of which are reasoning models) on GRAB-Lite and report the results in Tab. 6. The original baselines attain similar overall scores on both GRAB and GRAB-Lite, with Claude 3.5 Sonnet scoring highest on both. The frontier models also perform well, with GPT-5 achieving the highest overall score, approaching 50%. An interesting observation is the strong performance of the newly evaluated models on the Transforms task, which was the lowest scoring for the GRAB baselines. This suggests the crossing of a capability threshold where the models can now perform the abstract graph transformations and interpret the resulting function to a reasonable degree of accuracy. o1 averaged >>5k completion tokens (\$0.33) per question, highlighting the need for such a lightweight eval.
5 Conclusions
We introduce GRAB, a GRaph Analysis Benchmark for LMMs featuring five tasks covering a diverse range of graph configurations and settings. We leverage synthetic data generation to create high-quality, difficult questions that prove challenging for frontier LMMs, with the best-performing model attaining a score of 21.0%. On GRAB, we evaluate 20 closed and open-source models, attaining a broad profile of the abilities of the current generation of LMMs. During inference, we use lightweight prompts with clear output format instructions and assess answer correctness by exact string matching. This approach jointly assesses the abilities of models to perform both the visual perception and reasoning in our tasks and reliably follow output instructions. We find little difference in performance using exact matching evaluation vs. automatic LLM evaluation, suggesting most current models can sufficiently follow output instructions, validating exact matching as a suitable evaluation protocol. We also curate a lightweight GRAB-Lite question set designed to evaluate reasoning models. Finally, we conduct studies revealing the categories and splits the models struggle with, and the impacts of different question types and plotting libraries. We release our datasets along with generation and evaluation code for the community to use and hope our work encourages research towards the next generations of LMMs.
Acknowledgements
This work was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (reference EP/S022961/1), an Isaac Newton Trust grant, a research gift from Google, an EPSRC HPC grant, the Hong Kong Research Grant Council - General Research Fund (Grant No. 17211024), and HKU Seed Fund for Basic Research. Samuel would like to acknowledge the support of Z. Novak and N. Novak in enabling his contribution.
Appendix
We structure this Appendix into 9 parts. In the first part (§A), we include curation details for the Screenshots and Noise splits of the Real task questions. In the next four parts, we report additional results and ablations, including a question format ablation (§B), a comparison of different performance metrics (§C), model performance results on an OCR-based graph analysis question set (§D), and a graph plotting libraries performance ablation (§E) and display example plots with different libraries. Finally, we detail the inference settings used for GRAB-Lite evaluation (§F), include examples of the prompts used for inference and LLM evaluation (§G) along with details of the compute costs of our work (§H) and the specific API model versions used throughout this evaluation (§I).
Appendix A Real task curation details
A.1 Screenshots
To construct the Screenshots split, after creating the question graphs, we randomly selected one of the following digital contexts to embed the image:
  • Presentations (Google Slides, PowerPoint)
  • Video Calling (Teams, Zoom, Meet)
  • Image Viewing (Preview)
  • IDE (VSCode)
  • Email
  • Word Processing (Word, Google Docs)
We curated a set of different designs for each digital context (screenshot), including different combinations of background applications/OS. To each, we added an arbitrarily coloured box to the screenshots, denoting where the graph would be embedded. Once created, the coloured box was then replaced with the graph and the composite figure saved and used as part of the question.
A.2 Noise
To construct the Noise split, after initially creating the graphs, we randomly applied one of the following types of noise to the image:
  • Gaussian noise
  • Salt and pepper noise
  • Brightness/contrast
  • Blur
  • Spatter/smear
  • JPEG artifacts
  • Rotations
  • Flips
Appendix B Question format ablation
@@
Appendix C Additional metrics
For a broader view, we report some additional metrics in Tab. 8. In addition to the accuracy scores reported in the main paper (pass@1), here we also report pass@5, 5/5 reliability, root mean squared error (RMSE) and mean absolute error (MAE). These additional metrics largely preserve the pass@1 model rankings, though there is some variation. Our core analysis focuses on pass@1 accuracy as each alternative metric has limitations, making use on GRAB unfeasible. While the pass@5 and 5/5 reliability metrics provide insights into model performance outside of near-deterministic settings, running each evaluation on 3284 questions five times is impractical. Although most GRAB answers are numeric, distance-based error metrics, such as RMSE and MAE are problematic due to differing scales, ground-truth equal to 0 and treating non-numeric LMM answers.
Performance metric
pass@1 $\uparrow$ pass@5 $\uparrow$ 5/5 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
Claude 3.5 Sonnet 18.6 22.7 14.0 297.5 29.7
Gemini 1.5 Pro 21.9 28.9 14.7 209.9 28.3
GPT-4o 18.0 26.0 8.3 226.9 27.7
Table 8: Performance metrics on Real Screenshots split.
Appendix D OCR experiments
Model Accuracy (%)
Random chance 20.0
Closed-source LMMs
Claude 3 Haiku [3] 42.9
Claude 3 Sonnet [3] 91.4
Gemini 1.0 Pro Vision [13] 100.0
Gemini 1.5 Flash [49] 100.0
Gemini 1.5 Pro [49] 100.0
GPT-4 Turbo [36] 97.1
GPT-4V [36] 97.1
GPT-4o [37] 82.9
Reka Core [46] 97.1
Reka Edge [46] 88.6
Reka Flash [46] 100.0
Open-source LMMs
TransCore-M [48] 80.0
Yi-VL-6b [1] 51.4
OmniLMM-3b [44] 65.7
Qwen-VL-Chat [8] 71.4
Table 9: Accuracy scores on OCR experiments.
Appendix E Plotting libraries ablation
To compare different plotting libraries, we sample 100 GRAB questions and synthesise them using both Matplotlib [22] and the Seaborn [56] library to create two sets of questions that differ in the aesthetics and styles of the libraries but are otherwise identical. The near agreement of Claude 3.5 Sonnet’s scores on these sets (Tab. 10) suggests that plotting with different libraries does not significantly impact overall performance; therefore, we focus on Matplotlib plots for GRAB. However, to increase the diversity of GRAB-Lite, we re-generate half of the synthetic plots using Seaborn. Fig. 7 displays examples of identical functions and series plotted with both the Matplotlib and Seaborn libraries.
Accuracy (%)
Properties Functions Series Transforms Overall
Plotting Library (25) (25) (25) (25) (100)
Matplotlib [22] 48.0 16.0 36.0 20.0 30.0
Seaborn [56] 40.0 20.0 24.0 20.0 26.0
Table 10: Claude 3.5 Sonnet accuracy on different plotting libraries on a 100-question subset of GRAB.
Appendix F GRAB-Lite inference
For inference on the GRAB-Lite questions we evaluated o1, GPT4.1 and the GPT-5 models via the OpenAI API [40]. The Gemini Developer API [20] was used to evaluate Gemini 2.0 Flash and Flash Lite, Gemini 2.5 Flash and Flash Lite, and Gemini 2.5 Pro. We used the Vertex AI API [14] to evaluate Claude 3.7 Sonnet and Claude Sonnet 4 and 4.5, and the xAI API [59] for Grok 4. Greedy decoding was used where possible. The ‘reasoning effort’ parameter for o1, GPT-5 mini and nano was set to ‘high’; for GPT-5 it was set to ‘medium’.
Appendix G Prompts
@@
Appendix H Compute
As we restricted the size of our benchmark to 3284, the total compute required for this work is relatively low. Negligible compute was needed to create the synthetic images. Inference with the closed-source models was carried out via API calls. Using a single NVIDIA A100-80GB GPU, inference on the entire GRAB benchmark using LLaVA-1.5 7b can be carried out in approximately 45 minutes, using a single process (inference time is significantly reduced with multiprocessing).
Appendix I API model versions
@@
  • Gemini 2.0 Flash:
    gemini-2.0-flash-001
  • Gemini 2.0 Flash Lite:
    gemini-2.0-flash-lite-001
  • Gemini 2.5 Flash:
    gemini-2.5-flash-preview-05-20
  • Gemini 2.5 Flash Lite:
    gemini-2.5-flash-lite
  • Gemini 2.5 Pro:
    gemini-2.5-pro
@@
  • Claude 3.7 Sonnet:
    claude-3-7-sonnet@20250219
  • Claude Sonnet 4:
    claude-sonnet-4@20250514
  • Claude Sonnet 4.5:
    claude-sonnet-4-5@20250929
@@
  • Grok-4:
    grok-4-0709
(a) Matplotlib Functions example
(b) Seaborn Functions example
(c) Matplotlib Series example
(d) Seaborn Series example
Figure 7: Example equivalent plots from the Matplotlib and Seaborn plotting libraries.